OpenStudy (anonymous):

8th grade math!!! I will give metals!!!!

OpenStudy (misty1212):

are they precious metals?

OpenStudy (misty1212):

one at a time dear

OpenStudy (anonymous):

(07.02 MC) An equation is shown below: 4(2x – 3) = 1 Which of the following correctly shows the first two steps to solve this equation? (1 point) Step 1: 6x – 7 = 1; Step 2: 6x = 8 Step 1: 6x + 1 = 1; Step 2: 6x = 0 Step 1: 8x – 12 = 1; Step 2: 8x = 13 Step 1: 8x – 3 = 1; Step 2: 8x = 4

OpenStudy (anonymous):

thats alot

OpenStudy (anonymous):

of work

OpenStudy (anonymous):

I just only posted 1 question you might want to read before you post